OpenStudy (anonymous):

Give an example of a function, f(x), that has an inflection point at (1,4)

OpenStudy (anonymous):

y-4=(x-1)^3

OpenStudy (anonymous):

x^3-1 maybe.. i dont have a graphing calculator available though so i cant check it

OpenStudy (anonymous):

i.e. f(x)=4+(x-1)^3

OpenStudy (anonymous):

that makes more sense. the wording has been confusing me lol. would appreciate if you could help on this one too.

OpenStudy (anonymous):

An example of a function, g(x), that has a local maximum at (-3,3) and a local minimum at (3,-3)

OpenStudy (anonymous):

f(x)=-3sin(x*pi/6)

OpenStudy (anonymous):

Thanks :)